Data below is gathered:

ROE = 14%

Sales = $10,000

Total Debt Ratio = .25

Total Liabilities = $100,000

What is ROA?

A. 10.50%

B. 56%

C. 3.50%

D. 18.67%

E.   14%
Answer #1

Answer: The correct answer is A i.e. 10.50%

Equity Ratio = 1-Debt ratio
Equity Ratio = 1-0.25
Equity Ratio = 0.75

Equity Multiplier = 1/Equity Ratio
Equity Multiplier = 1/0.75
Equity Multiplier = 1.3333

ROE = Equity multiplier * ROA
0.14 = 1.33 * ROA
ROA = 0.14 /1.3333
ROA = 0.1050 or 10.50